OpenStudy (anonymous):

Write the paranetric functions x=tant and y=sect in form y=f(x) by eliminating t

OpenStudy (gorv):

we know that sec^2t - tan^2t=1 x^2=tan^2t y^2=sec^2t y^2-x^2= sec^2t - tan^2t y^2-x^2=1 y^2=1+x^2 y=(1+x^2)^(1/2)

OpenStudy (anonymous):

\[1 + \tan^{2} t= \sec^{2} t\] \[1 + x^{2} = y^{2} \] \[\implies y = \sqrt{1 + x^{2}}\]